Operations Report – April 30th

Crew 193 Operations Report 30Apr2018

SOL: 8

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10. Soldering Irons

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now). At 4:45 PM today, the SoC was 73%, which is lower than it should be, given how sunny it is today; there are some storm clouds to the southwest, which could be blocking the sun, we need to keep an eye on SoC.
I will repair charging plugs once we have a soldering iron.

Generator (hours run): 10:45 pm to 8:10 am.

Solar— SOC 59 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0.5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134370.9

Toilet tank emptied: no
Date last emptied: April 29

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 300, 350.1, 350.3

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 2 Gallons

# Hours the ATVs were Used today: 3

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.

Summary of suits and radios: Still work to do on two piece suits. Repair of suit chargers still in progress. Other than #10, one piece suits nominal.

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment proceeding.

Summary of RAM operations: A charging connector was moved from RAM for space suit repair (once soldering iron arrives)

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: We would like some black straps used for connecting two piece suit neck rings to backpacks. I did not see any in the RAM.

Copyright © The Mars Society. All rights reserved. | Main Site